OpenStudy (anonymous):

what is the value for x in this equation? log(6x+5)-log2=log5 a) -0.833 b)-0.667 c) 0.333 d) 0.667 e) 0.833

Nnesha (nnesha):

\[\huge\rm \log(6x+5)-\log (2)=\log(5)\] like this ?

Nnesha (nnesha):

familiar with log rules ?

OpenStudy (anonymous):

yes like that, and no not familiar

Nnesha (nnesha):

`quotient rule`\[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division `product rule` \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication `power rule` \[\large\rm log_b x^y = y \log_b x\] \[\rm \log x= \log y \rightarrow x=y\]

Nnesha (nnesha):

there is a negative sign so apply quotient rule \[\large \rm \color{Red}{ \log(6x+5)-\log (2)}=\log(5)\]

Nnesha (nnesha):

here is an example \[\log2 - \log 3 \rightarrow \log \frac{2}{3} \]